OpenStudy (anonymous):

Find the points of intersection of r=3sinθ and r=1+sinθ.

OpenStudy (amistre64):

 i got no idea if thise are to scale

OpenStudy (amistre64):

1+sin(t) might be more of a limcon tho

OpenStudy (amistre64):

r=3sinθ and r=1+sinθ 3sinθ =1+sinθ 3sinθ -sinθ = 1 (3-1)sinθ = 1 sinθ = 1/2

OpenStudy (anonymous):

Thank you! That is exactly what I needed! :D

OpenStudy (amistre64):

yw